OpenStudy (anonymous):

find a quadratic equation with the roots -1+4i and -1-4i

OpenStudy (jango_in_dtown):

quadratic equation with roots x=a and x=b are given by (x-a)(x-b)=0

OpenStudy (jango_in_dtown):

so we have (x-(-1+4i))(x-(-1-4i))=0

OpenStudy (jango_in_dtown):

implies (x+1-4i)(x+1+4i)=0 implies (x+1)^2+16=0

OpenStudy (jango_in_dtown):

implies x^2+2x+17=0

OpenStudy (jango_in_dtown):

@Hope214

OpenStudy (anonymous):

i dont understand can you please explain?

OpenStudy (jango_in_dtown):

ok.. see if the roots are a and b, the equation is given by (x-a)(x-b)=0

OpenStudy (jango_in_dtown):

here a=-1+4i and b=-1-4i so plug them in the equation (x-a)(x-b)=0

OpenStudy (anonymous):

ok so once you plug them in what do you do

OpenStudy (jango_in_dtown):

then we reduce it in the general quadratic form which is px^2+qx+r=0 where p,q,r are constants

OpenStudy (jango_in_dtown):

now lets get back to the problem, after substitutiong we have (x-(-1+4i))(x-(-1-4i))=0 implies (x+1-4i)(x+1+4i)=0

OpenStudy (jango_in_dtown):

implies {(x+1)-4i}{(x+1)+4i}=0 now use the formula (a-b)(a+b)=a^2-b^2 and you get (x+1)^2-(4i)^2=0 implies (x+1)^2+16=0

OpenStudy (jango_in_dtown):

implies x^2+2x+1+16=0 implies x^2+2x+17=0

OpenStudy (anonymous):

ok thank u